### v2

@@ 6.2 System Benchmarking
6.3 Correctness Evaluation
Figure 10 shows the results of the correctness of RAG evaluation, presenting both the overall results and assessments based on predefined categories. The results demonstrate remarkable consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents’ navigation tasks. The slight percentage of incorrect answers remains within an acceptable threshold, given the specific context employed and the inclusion of general questions.
@@ 6.3 Correctness Evaluation
Regarding individual categories, the correctness in the category ’Navigation Process Overview’ is relatively consistent across all scenarios, indicating a good understanding of events and progress related to navigation. However, the broad nature of the question "What has happened in this ROS 2 log regarding navigation?" initially yielded a correctness below 65% in the third scenario. In this instance, the cancellation of a goal resulted in the execution and failures of Nav2 Behavior Tree nodes, not included in the other two scenarios and focused on navigation recovery. Therefore, the incorporation in a dataset of information related to the cancellation of a goal results in LLM answers that usually diverge from the ground truth provided, as the model occasionally prioritized Behavior Trees information, other times focused on locations or timestamps. This variability made it difficult to align with the ground truth provided, even if the obtained answers could be considered partially accurate. Introducing the prompt extension "Describe the sequence of events regarding navigation goals and include information about each of them." in question 1 significantly increased correctness in the first category, achieving results around and exceeding 75% in all three scenarios. This fact underscores the effectiveness of prompt engineering when tackling ambiguous or broad queries.
@@
Through questions included in the Nav2 Behavior Tree and node status category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs [80]. While the model excels in identifying Behavior Tree Nodes given its description, it faces challenges in pinpointing these components under specific conditions. As a result, the category correctness suffers a drop in Scenario 3, although overall results remain above 75%. In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 Behavior Tree nodes not present in the other scenarios. Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 Behavior Tree node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation. This can be attributed to the fact that the complexities of the recovery actions and the dynamic nature of Behavior Tree executions during navigation recovery introduce complexities that the model may not fully generalize from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.
6.4 Criteria-based Evaluation
Figure 11 shows the remarkable performance of the selected LLM in the evaluated metrics, demonstrating consistency with slight variations in the scores. On average, the answers provided are accurate and in line with the reference, though there might be room for improvement. While responses generally provide useful information, there are instances where additional depth could enhance their utility. In addition, most answers usually include a remarkable balance between conciseness and effectiveness, showing a proficient level of brevity, as well as clear and logical organization. The overall results show no notable differences in these metrics across the three scenarios. This suggests that messages within the Rosbag related to the introduction of obstacles or cancellation of goals do not notably influence the general criteria evaluation outcomes.
@@ 6.4 Criteria-based Evaluation
Concerning general questions about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3. While coherence and conciseness are even improved, Scenario 1, involving direct navigation to three goals, outperforms in accuracy and helpfulness. This fact suggests that effective communication on general queries benefits from the absence of obstacles or goal cancellations, enhancing overall understanding.
@@
In the category ’Trajectory Planning and Re-planning’, the questions are effectively addressed in all three scenarios, showing scores consistently above 8 in every evaluated metric, even when obstacles are introduced or a goal is cancelled. These results underline the system’s proficiency in handling trajectory-related questions despite potential new challenges.
@@
Regarding the category ’Goal Completion and Navigation Task Status’, Scenarios 1 and 2 demonstrate the effectiveness of the system in conveying goal completion and navigation task status information. Nevertheless, the slightly lower scores for Scenario 3 indicate potential challenges in effectively communicating completion status when a goal is aborted.
@@
Concerning the ’Specifics about Goals and Locations’ category, the system consistently provides valuable goal and location information, even in the presence of obstacles or goal cancellations. This suggests a robust capacity to understand and localize specific details regardless of the challenges introduced in each scenario.
@@
The results regarding questions included in ’Nav2 Behavior Tree and Node Status’ demonstrate effectiveness in communication across all scenarios and metrics. This emphasizes the potential for Behavior Trees to complement language models in explainability, offering insight into the decision-making processes during navigation tasks.
In summary, although results can be improved in accuracy regarding general questions, the language model exhibits adaptability to different navigation scenarios, facing obstacles and changes in navigation tasks. The consistency in conveying information related to Nav2 Behavior Trees suggests a robust understanding of underlying navigation processes. The model also performs well when providing information about trajectory planning, goal completion, and specifics about goals and locations, regardless of the specific challenges introduced in each scenario. Its robustness in effective communication, despite technical vocabulary, details and the introduced challenges, is demonstrated through the achievement of responses that are not only generally accurate but also easily comprehensible to users without expertise, which enhances the model’s practical utility.
@@ 7 Conclusions
This work presents an accountability and explicability architecture, implemented for Robot Operating System (ROS)-based mobile robots. Our approach comprises a black box component for accountability that effectively addresses real-time recording challenges, ensuring message integrity and verifiability through blockchain technology, and an explainability component. This last component is based on the use of Retrieval Augmented Generation (RAG) and Large Language Models (LLMs) through a private question-answering system designed to use the information recorded by the first component as context. This solution demonstrates the feasibility of obtaining accurate, coherent, concise and helpful explanations for robot actions from raw data stored in Rosbag files, thereby enhancing human-robot interaction. Additionally, both main components have been integrated through the development of a third main functionality in charge of filtering non-relevant information and making messages closer to the final user, due to their specific nature can introduce challenges in their comprehension process. Furthermore, this last component releases software developers from making direct calls to the logging Application Programming Interface (API), easing the development process and improving the scalability and adaptability of our solution for every navigation task performed by a ROS-based robot.
@@
While our approach has shown remarkable results in navigation functionalities regarding several scenarios and circumstances, future work should involve adapting our solution to other applications by simply adapting the recording and the processing components to the relevant topics of the area or domain of use. Furthermore, the consistency and robustness of responses generated by our explainability approach could benefit from refinement across different scenarios, especially when categorizing questions under specific subjects. Therefore, future work should explore the impact of In-Context Learning (ICL) techniques on prompt generation, in addition to fine-tuning the model for the specific context of autonomous agent navigation tasks.
@@ A.2 Explanations Obtained
The 288 questions and answers obtained through the explanation generation and evaluation process are available on the aforementioned mentioned Github repository  [^4] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
@@
. Additionally, a Python notebook has been developed to interact with the content of the previous sheet. This notebook is available on  [^5] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
@@ Acknowledgments
This research has been partially funded by the Recovery, Transformation, and Resilience Plan, financed by the European Union (Next Generation) thanks to the TESCAC project (Traceability and Explainability in Autonomous Systems for improved Cybersecurity) granted by INCIBE to the University of León; and by grant PID2021-126592OB-C21 funded by MCIN/AEI/10.13039/501100011033 EDMAR (Explainable Decision Making in Autonomous Robots) project, PID2021-126592OB-C21 funded by
MCIN/AEI/10.13039/501100011033 and by ERDF ”A way of making Europe”.

### v3

@@ 6.2 System Benchmarking
6.3 Correctness Evaluation Using LLM-as-a-judge
Figure 10 shows the results of the correctness of RAG evaluation, presenting both the overall results and assessments based on predefined categories. The results demonstrate consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents’ navigation tasks. The slight percentage of incorrect answers remains within an acceptable threshold, given the specific context employed and the inclusion of general questions.
@@ 6.3 Correctness Evaluation Using LLM-as-a-judge
Regarding individual categories, the correctness in the category ‘Navigation Process Overview’ is relatively consistent across all scenarios, indicating a good understanding of events and progress related to navigation. However, the broad nature of the question "What has happened in this ROS 2 log regarding navigation?" initially yielded a correctness below 65% in the third scenario. In this instance, the cancellation of a goal resulted in the execution and failures of Nav2 Behavior Tree nodes, not included in the other two scenarios and focused on navigation recovery. Therefore, the incorporation in a dataset of information related to the cancellation of a goal results in LLM answers that usually diverge from the ground truth provided, as the model occasionally prioritized BTs information, other times focused on locations or timestamps. This variability made it difficult to align with the ground truth provided, even if the obtained answers could be considered partially accurate. Introducing the prompt extension "Describe the sequence of events regarding navigation goals and include information about each of them." in Question 1 significantly increased correctness in the first category, achieving results around and exceeding 75% in all three scenarios. This fact underscores the effectiveness of prompt engineering when tackling ambiguous or broad queries.
@@
Through questions included in the ‘Nav2 Behavior Tree and Node Status’ category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs [110]. While the model excels in identifying BT nodes given its description, it faces challenges in pinpointing these components under specific conditions. As a result, the criteria correctness suffers a drop in Scenario 3, although overall results remain above 75%. In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 BT nodes not present in the other scenarios. Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 BT node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation. This can be attributed to the fact that the complexities of the recovery actions and the dynamic nature of BT executions during navigation recovery introduce complexities that the model may not fully generalize from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.
6.4 Criteria-based Evaluation Using LLM-as-a-judge
Figure 11 shows the performance of the selected LLM in the evaluated metrics, demonstrating consistency with slight variations in the scores. On average, the answers provided are accurate and in line with the reference, though there might be room for improvement. While responses generally provide useful information, there are instances where additional depth could enhance their utility. In addition, most answers usually include a good balance between conciseness and effectiveness, with clear and logical organization. The overall results show no notable differences in these metrics across the three scenarios. This suggests that messages within the Rosbag related to the introduction of obstacles or cancellation of goals do not notably influence the general criteria evaluation outcomes.
@@ 6.4 Criteria-based Evaluation Using LLM-as-a-judge
Concerning general questions about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3. While coherence and conciseness improve, Scenario 1, which involves direct navigation to three goals, scores higher in accuracy and helpfulness. This fact suggests that effective communication on general queries benefits from the absence of obstacles or goal cancellations, enhancing overall understanding.
@@
In the category ‘Trajectory Planning and Re-planning’, questions are effectively addressed across all three scenarios, showing scores above 8 in every metric, even when obstacles or goal cancellations are introduced. These results highlight the system’s capability to manage trajectory-related queries effectively despite challenges.
@@
Regarding the category ‘Goal Completion and Navigation Task Status’, Scenarios 1 and 2 effectively convey information about goal completion and navigation task progress. Nevertheless, slightly lower scores in Scenario 3 indicate challenges in effectively communicating completion status when a goal is aborted.
@@
In the ‘Specifics about Goals and Locations’ category, the system provides valuable goal and location information, even in the presence of obstacles or goal cancellations. This suggests a strong ability to understand and localize specific details regardless of the challenges introduced in each scenario.
@@
The results regarding questions included in ‘Nav2 Behavior Tree and Node Status’ category demonstrate effectiveness in communication across all scenarios and metrics. This emphasizes the potential for BTs to complement language models in explainability, offering insight into the decision-making processes during navigation tasks.
In summary, although results can be improved in accuracy for general questions, the language model shows adaptability across different navigation scenarios, including those with obstacles and task changes. The consistency in conveying information related to Nav2 BTs suggests a solid understanding of underlying navigation processes. The model also performs well when providing information about trajectory planning, goal completion, and specifics about goals and locations, regardless of the specific challenges introduced in each scenario. Nevertheless, while evaluation using LLMs excels in scalability and efficient resource usage, it is still recommended to test a subset of questions with human evaluators. This ensures a deeper understanding of how well the approach aligns with human expectations and effectively communicates, particularly in scenarios involving complex or nuanced details.
6.5 Human Evaluation of LLM-generated Explanations
This section presents an analysis of the results from the human evaluation conducted with 17 participants, as stated in Section 5.2.2.
6.5.1 Distribution Analysis
Figure 13 depicts the distribution of Likert scale responses for the five evaluation criteria across the three scenarios, showing the frequency of each response value. The highest ratings predominate across all scenarios and criteria, suggesting that participants generally found the explanations satisfactory. However, Scenario 2 and Scenario 3 exhibit greater variability, with an increase in mid-range scores for certain criteria, such as Usefulness. These trends are particularly noticeable in broad questions, such as those in Category 1 regarding the navigation process overview, or when the model’s answers do not align with the ground truth. In such cases, participant opinions diverge significantly, even in the presence of factual inaccuracies, indicating a potential bias in how the utility is perceived when discrepancies with the ground truth arise.
Figure 13: Distribution of Likert scale responses across scenarios and evaluation criteria.
6.5.2 Reliability Analysis
The reliability analysis resulted in an overall Cronbach’s Alpha value of 0.9792, indicating good internal consistency and confirming that the evaluation criteria reliably capture the intended dimensions of the robot’s explanations.
Additionally, we calculated Cronbach’s Alpha if an item is deleted for each criterion. This metric helps assess the impact of removing an individual item on the scale’s overall reliability. As summarized in Table 5, the "Alpha if Deleted" values are consistently close to the overall Cronbach’s Alpha. Specifically, these values range from 0.9699 (Usefulness) to 0.9798 (Informativeness), with all values slightly lower than the overall result. This suggests that removing any single item would only have a minimal impact on the reliability of the scale, indicating that each criterion contributes positively to the measurement of overall quality.
Further, we examined the item-total correlations, which quantify the strength of the relationship between each criterion and the total score. These correlations range from 0.9286 for the Informativeness criteria to 0.9660 for Usefulness. The high item-total correlations support the relevance of each item in assessing the key dimensions of the robot’s explanations.
Table 5: Cronbach’s Alpha and item-wise "Alpha if Deleted" in human evaluation.
Criteria Cronbach’s Alpha if Deleted Item-Total Correlation
Completeness 0.9724 0.9492
Informativeness 0.9798 0.9286
Satisfaction 0.9752 0.9567
Understandability 0.9724 0.9502
Usefulness 0.9699 0.9660
Finally, we assessed the inter-item correlations, which provide insight into how strongly the criteria correlate with one another. The inter-item correlation matrix, presented in Table 6, shows values ranging from 0.878 (between Informativeness and Satisfaction) to 0.981 (between Satisfaction and Usefulness). These high correlations indicate that the criteria are closely related but not redundant, as each captures a unique and complementary aspect of the evaluation process.
Table 6: Inter-Item correlation matrix in human evaluation.
Completeness Informativeness Satisfaction Understandability Usefulness
Completeness 1.000 0.976 0.905 0.911 0.911
Informativeness 0.976 1.000 0.878 0.887 0.881
Satisfaction 0.905 0.878 1.000 0.936 0.981
Understandability 0.911 0.887 0.936 1.000 0.947
Usefulness 0.911 0.881 0.981 0.947 1.000
6.5.3 Correlation Analysis
Building on the findings from the reliability analysis, we further explored the relationships among the variables in our dataset through Pearson correlation matrices. While Cronbach’s Alpha results demonstrated the internal consistency of the evaluation criteria, the correlation analysis extends this by examining how these criteria relate to each other across different scenarios and explanation categories.
Figure 14 presents the Pearson correlation matrix for participants’ responses across the three scenarios. It reveals strong associations, indicating a high level of consistency in participants’ assessments. This consistency suggests that the explanations provided were clear and easily comparable, allowing participants to reliably evaluate them regardless of the specific context of the robot’s task.
Figures 15 and 16 depict the relationships among evaluation criteria and question categories. Regarding the criteria, Completeness and Informativeness exhibit a strong positive correlation, as do Understandability and Usefulness. This indicates that participants are more likely to perceive explanations as useful when they are also clear. In contrast, the correlation between Informativeness and Usefulness is weaker (0.55), highlighting a divergence in participants’ evaluations. This divergence may stem from biases previously discussed, where participants find it challenging to judge the usefulness of an explanation that is not fully accurate.
Regarding the question categories, the ‘Navigation Process Overview’ and ‘Specifics About Goals and Locations’ categories exhibit a strong correlation (0.99), indicating that participants tended to provide similar ratings for explanations in these areas. This suggests that these types of explanations are perceived similarly. In contrast, the ‘Trajectory Planning and Re-planning’ category shows weak or negative correlations with other categories. This discrepancy arises from differences in evaluating broad, open-ended explanations versus more specific ones.
Figure 14: Correlation matrix of responses by scenario.
Figure 15: Correlation matrix of responses by criteria.
Figure 16: Correlation matrix of responses by category.
6.5.4 Inferential Statistics
Statistical analysis often requires understanding the underlying data distribution to ensure the appropriate methods are applied. Since parametric tests assume normality, violations of this assumption can compromise the validity of the results. To assess normality, we applied the Shapiro-Wilk test, a robust method for small to moderate-sized samples, across various groupings, including scenarios, categories, questions, and criteria.
The test results consistently revealed significant deviations from normality, with $p$-values well below 0.05, leading to the rejection of the null hypothesis of normality for all groups. These findings confirm that the dataset does not satisfy the assumptions required for parametric tests, such as ANOVA or t-tests. As a result, non-parametric methods were adopted in subsequent analyses to ensure the robustness and validity of statistical inferences.
The Friedman test was used due to its ability to assess differences in related groups when the data does not meet the assumptions of normality required for parametric tests. The results of the Friedman test across the three scenarios reveal significant differences in participant ratings for several explanations, indicating varying levels of agreement across conditions. In Scenario 1, significant differences were observed for explanations to Questions 1 and 2, suggesting diverse responses. In contrast, most other questions in this scenario showed no significant differences ($p\geq 0.05$).
Explanations 1 and 2 in all three scenarios addressed broad questions, whose answers diverged from the ground truth, as discussed in the GPT-4 evaluation in 6.3. Broad questions often result in responses that prioritize varied information, such as BTs, locations, or timestamps, making it difficult to align the answers with the ground truth. This fact justifies the significant variability in participant ratings for these explanations.
In Scenario 2, significant differences were observed for Explanations 1, 7, 8, and 15, with $p$-values below the 0.05 threshold, while other questions did not show significant variations. Specifically, the explanations for Questions 7 and 8 were misaligned with the ground truth, leading to significant variability in participant assessments. In these cases, participants’ ratings lacked a clear trend across the evaluated criteria. A similar pattern was observed in Explanation 15 of Scenario 2, where the model provided irrelevant information that did not directly address the user’s inquiry. This lack of relevance and alignment likely contributed to the observed variability in participant ratings.
Finally, Scenario 3 showed significant differences for Explanations 1, 2, 7, and 16, indicating that the conditions for these explanations led to notable variations in participant ratings. Explanation 7, does not address the answer to the corresponding question, including no relevant information. Explanation 16 in Scenario 3 caused variability in participants’ answers due to its incomplete nature, highlighting the importance of content clarity in generating consistent and meaningful explanations. These findings suggest that while most explanations remain consistent, certain ones resonate differently depending on the scenario, nature of the question, and LLMs indeterminism.
Figure 17: Friedman test results.
The post-hoc Nemenyi test was conducted for explanations that showed significant differences in participant ratings based on the Friedman test results. This allowed us to identify specific participant pairs with notable divergences in their responses across the three scenarios. The Nemenyi test revealed that most participants were consistent in their evaluations, but certain pairs exhibited significant variability. For example, in Scenario 1, the evaluation for Explanations 1 and 2 highlighted differences involving Participant 8 and several others. In Scenario 2, significant differences emerged for Explanation 1 among Participants 3, 6, 7, and 13. Similarly, in Scenario 3, variability was observed for Explanations 1 and 16, where selected participants differed significantly from others. These discrepancies are likely due to natural human biases, differences in interpretation, and subjective judgments that are inherent in human responses. Overall, while the explanations were generally clear and consistent for most participants, the observed variability underscores the role of individual perspectives and biases in shaping evaluations, reflecting the inherent complexity of human judgment.
6.6 Similarity Between Human and LLM Evaluations
As described in 5.2, previous studies have demonstrated that LLMs can effectively approximate human judgment in various evaluation tasks, including those involving subjective interpretation. In this work, we conducted a focused validation of LLM-generated evaluations by comparing a subset of 48 explanations from one of six evaluation runs performed by the LLM against human assessments. This subset included 16 explanations per scenario, allowing us to validate the LLM’s judgment within a practical scope.
Figures 18 and 19 depict the mean responses from 17 human participants and GPT-4, respectively, acting as a judge in this subset. The comparison highlights notable similarities in scoring patterns across scenarios and explanations, suggesting a high degree of alignment between human and LLM evaluations. Both evaluations consistently identify high-quality explanations with scores clustered around 4.5–5, indicating shared agreement on the quality of certain explanations. Similarly, lower-rated explanations identified by humans are also flagged by the LLM, revealing a mutual recognition of weaker content.
However, while the overall trends are similar, there are some differences in the exact ratings and outlier values between the two evaluations. These differences could be attributed to nuances in how the LLM interprets the explanations compared to human evaluators, who might take into account subtle contextual or subjective factors that an LLM might miss. Despite these slight variations, the visual correspondence between the human and LLM heatmaps indicates that the LLM provides a reliable approximation of human judgment, making it a promising tool for scaling evaluation processes without sacrificing much in terms of accuracy. This similarity could be valuable in demonstrating that LLMs are suitable alternatives for tasks like explanation evaluation, especially in large-scale studies where human evaluation may be impractical.
Figure 18: Heatmap of mean responses from human evaluators Across Scenarios and Explanations.
Figure 19: Heatmap of mean responses from the LLM-as-Judge evaluation across scenarios and explanations.
@@ 7 Conclusions
This work presents an accountability and explicability architecture, implemented for Robot Operating System (ROS)-based mobile robots. Our approach comprises a black box component for accountability that effectively addresses real-time recording challenges, ensuring message integrity and verifiability through blockchain technology, and an explainability component. This last component is based on the use of Retrieval Augmented Generation (RAG) and Large Language Models (LLMs) through a private question-answering system designed to use the information recorded by the first component as context. This solution demonstrates the feasibility of obtaining accurate, coherent, concise and helpful explanations for robot actions from raw data stored in Rosbag files, providing satisfactory results for users and thereby enhancing human-robot interaction. Additionally, both main components have been integrated through the development of a third main functionality in charge of filtering non-relevant information and making messages closer to the final user, due to their specific nature can introduce challenges in their comprehension process. Furthermore, this last component releases software developers from making direct calls to the logging Application Programming Interface (API), easing the development process and improving the scalability and adaptability of our solution for every navigation task performed by a ROS-based robot.
@@
While our approach has shown promising results in navigation functionalities regarding several scenarios and circumstances, future work should involve adapting our solution to other applications by simply adapting the recording and the processing components to the relevant topics of the area or domain of use. Furthermore, the consistency and robustness of responses generated by our explainability approach could benefit from refinement across different scenarios, especially when categorizing questions under specific subjects. Therefore, future work should explore the impact of In-Context Learning techniques on prompt generation, in addition to fine-tuning the model for the specific context of autonomous agent navigation tasks.
@@ A.2 Explanations Obtained
The 288 questions and answers obtained through the explanation generation and evaluation process are available in the aforementioned mentioned Github repository  [^4] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
@@
. Additionally, a Python notebook has been developed to interact with the content of the previous sheet. This notebook at the following link  [^5] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
@@ Acknowledgments
This publication is part of the TESCAC project, financed “by European Union NextGeneration-EU, the Recovery Plan, Transformation and Resilience, through INCIBE”. In addition, this work has been partially funded by the project EDMAR, PID2021-126592OB-C21, funded by MCIN/AEI/10.13039/501100011033 and by ERDF A way of making Europe.